OpenStudy (hpfan101):

\[\lim_{x \rightarrow \infty} \sqrt{9x^2+x}-3x\]

OpenStudy (anonymous):

gimmick is to multiply by \[\frac{\sqrt{9x^2+x}+3x}{\sqrt{9x^2+x}+3x}\]

OpenStudy (hpfan101):

When I multiplied by the conjugate I got: \[\frac{ 9x^2+x-9x }{ \sqrt{9x^2+x}+3x }\]

OpenStudy (hpfan101):

Oh the denominator is supposed to be -x not +x

OpenStudy (hpfan101):

Actually nevermind, it's +x.

OpenStudy (anonymous):

\(-9x^{\color{red}2}\) in the numerator's second term

OpenStudy (hpfan101):

Oh ok.

OpenStudy (hpfan101):

Alright, I now know how to solve the question from there. Thank you guys!